# The LCLStream Ecosystem for Multi-Institutional Dataset Exploration

### v1

@@ 3.2. Data availability with LCLStream-API
LCLStream-API provides a web interface for external users to request data from LCLStreamer. It is built around LCLStreamer’s configuration file. As a REST-API, data transfers are started by POST-ing that configuration file as a typed JSON message to the transfers path. The response is either a validation error, or the ID for the newly created transfer. Issuing a GET or a DELETE to transfers/ID then reads the transfer status or stops a running transfer.
@@ 3.3. Message Buffering with NNG-Stream
Figure 3. NNG-Stream Connectivity diagram. Each cache stores messages from all producers in a circular buffer, and distributes them round-robin to all consumers in an at-most-once fashion. Connectivity is provided via NNG Push0/Pull0 socket types.(D’Amore, 2025) Multiple caches can work simultaneously to deliver traffic at rates on the order of tens of gigabytes per second.
Figure 3 shows the data flow pathway. Throughput tests run with a single cache on a laptop show aggregate bandwidth of 3 Gigabytes per second. These are limited only by local message routing and copying times. Network link speeds are quickly approaching the rate of 10s of Gigabytes per second. Hence, NNG-stream, if replicated to 3 or 4 simultaneous caches, is capable of saturating these network links.
@@ 3.4. Experiment Management with E-Log, ARP and Airflow
The ARP is flexible, and can start different types of workflows tied to run events. The two most common cases are simple shell scripts that submit jobs to a batch queue, or sophisitcated worflows managed by an Apache Airflow instance. Special Airflow operators have been written to interact with the S3DF cluster via the SLURM task management system, and with the Perlmutter supercomputer at NERSC via the Superfacility API (Enders et al., 2020). An additional operator that will allow communication with Psi-K (see 3.4) is currently under development, with the goal to allow Airflow to schedule tasks on a higher number of HPC facilities using different task scheduling systems. Within the context of the LCLStream project, we may use an Airflow operator to start an LCLStreamer-API transfer. This will start the data streaming pipeline as soon as a data collection run is started. Then compute-intensive analysis jobs will run remotely while the experiment is taking place.
@@ 3.5. HPC Job Management with Psi-K
The Psik and Psik-API projects provide an interface for batch queuing systems on local machines and HPC centers. As described above, Psi-k is organized around a file layout where a top-level directory stores a collection of jobs. Documents within a job are rigidly structured. Files contained within each job’s folder (jobs/JobID) include: the JobSpec, a record of job state changes, job logs, and working directory files (as a user-managed collection).
@@
This structure combines ideas from the ExaWorks Psi/J project(Hategan-Marandiuc et al., 2023) with traditional UNIX and cloud-based queuing systems. The working directory is optional. It was used here to avoid writing the configuration file as part of the job’s script.
Jobs are queued by POST-ing a JobSpec json to the jobs path. The server responds with either a validation error or a new JobID. Issuing a GET or a DELETE to jobs/JobID then gets information about the job or cancels a queued or running job. The result is is a single-document method for launching HPC jobs.
@@
Since each JobSpec names its backend, this extra information does not need to be a part of the API path (differing from NERSC’s Superfacility API(Enders et al., 2020)).
@@
Similar to Globus Compute,(Liu et al., 2021; Ananthakrishnan et al., 2024) the API listener can be run by individual users on a compute cluster’s login node. Alternatively, on Oak Ridge’s Frontier system, we run it on a Kubernetes pod with network access to the HPC cluster’s job scheduler and filesystem.(Papadimitriou et al., 2020) Differing from Globus compute, communications pass directly from the user to the Psik-API, bypasing a cloud services.
@@ 3.6. Mutual authentication with Certified
A companion signer package is designed to issue signed user certificates on a UNIX system. It can be deployed with essentially zero configuratoin by a typical HPC facility on any login node. Similar to MUNGE (the scheme for issuing user authentication credentials for launching jobs within the SLURM job scheduler it provides assertions about the user’s login name. In detail, it takes a certificate or a certificate signing request from a user, reads only the user’s public key, and issues the user a certificate linking their public key to their UNIX login name on that system. The user’s login name is determined by asking the kernel for the peer’s SO_PEERCRED information. Differing from MUNGE, it issues signed public keys. There is no danger of exposing these keys, since they can only be used by the original user who posesses the corresponding private key. Certified and signer never send the private key off of the user’s device.
@@ 4. Discussion
Discussions with the project team have highlighted the importance of continued, close collaboration between LCLS, S3DF and leadership compute facilities on these components. Although the team-developed microservices work in concert to accomplish streaming in their current form, greater integration with facility-deployed infrastructure will continue to improve their re-usability across sites and reliability for production work. Key areas of improvement include 1) eliminating manual intervention steps during install, startup, and execution (such as starting servers and creating ssh tunnels) 2) reducing site-specific adaptations needed to interoperate across different experiment and compute facility API-s, 3) standardizing security technologies (such as mutual TLS authentication), 4) solving the ”co-scheduling” problem for simultaneous commitment of beam time, network bandwidth, and compute resources, and 5) providing unencumbered access to fast, open-source data movement services (e.g. XRootD or FTS3) for managing site-to-site file copies.
@@ 5. Conclusion
This project has a unique position in relation to other efforts within the DOE Integrated Research Infrastructure (IRI) landscape.(Rogers, 2024) We have chosen to focus on small services that collaborate to provide a high-rate data channel. This avoids dealing with the full complexity of the experimental orchestration layer that Bluesky(Wijesinghe et al., 2024), INTERSECT(Brim et al., 2024), and GRETA/DELERIA’s(Cromaz et al., 2021) Janus framework do. At the same time, the certified, psik, and nng-stream packages have given general solutions to microservice communication, HPC job submission, and high-rate data buffering. They are targeted to simple, drop-in operations on HPC clusters, and avoid a centralized, cloud controller.
@@
Acknowledgments
We thank Daniel Pelfrey, Ross Miller and Jordan Webb for configuring high-speed network access to the Defiant system on OLCF’s ACE testbed. This research used resources of the SLAC National Accelerator Laboratory and the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. Use of the Linac Coherent Light Source (LCLS), SLAC National Accelerator Laboratory, is supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under Contract No. DE-AC02-76SF00515.

### v2

@@ 3.2. Data availability with LCLStream-API
LCLStream-API provides a web interface for external users to request data from LCLStreamer. It is built around LCLStreamer’s configuration file. As a REST-API, data transfers are started by POST operation, sending the configuration file as a typed JSON message to the transfers path. The response is either a validation error, or the ID for the newly created transfer. Issuing a GET or a DELETE to transfers/ID then reads the transfer status or stops a running transfer.
@@ 3.3. Message Buffering with NNG-Stream
Figure 3. NNG-Stream Connectivity diagram. Each cache stores messages from all producers in a circular buffer, and distributes them round-robin to all consumers in an at-most-once fashion. Connectivity is provided via NNG Push0/Pull0 socket types and, optionally ZeroMQ Push sockets.(D’Amore, 2025) Multiple caches can work simultaneously to deliver traffic at rates on the order of tens of gigabytes per second.
Figure 3 shows data flow through NNG-Stream. Throughput tests run with a single cache on a laptop show aggregate bandwidth of 3 Gigabytes per second. These are limited only by local message routing and copying times. Network link speeds are quickly approaching the rate of 10s of Gigabytes per second. Hence, NNG-Stream, if replicated to 3 or 4 simultaneous caches, is capable of saturating these network links.
In our setup, NNG-Stream is run on a data transfer node at S3DF. These nodes serve incoming connections from both compute nodes inside S3DF as well as compute nodes at HPC facilities such as OLCF and NERSC. Direct outbound traffic is technically possible for most compute facility setups, but is sometimes blocked administratively. Nevertheless, this traffic path must to be supported to create a useful HPC facility working environment. Working around administrative blocks like this is possible by doing things like running another NNG-Stream on a data transfer node or data streaming node at the HPC facility. However, setting up these extra hops adds complexity and cost while providing little added security benefit.
Failure properties of the data stream are inherited from the underlying Push/Pull socket model. Producers and consumers can connect and disconnect from the cache without impacting the streaming status. Hence, producer or consumer process failures will result in dropping the data residing in that process. However, MPI usually terminates when a node fails, which can cause a cascade failure. A cache failure will terminate the stream, requiring a restart.
Normal stream shutdown is triggered by sender disconnect events. When all senders have disconnected, the cache enters a drain state, where no new producer connections are allowed. When all its data has been sent, the cache disconnects and exits. Clients are setup to detect this disconnect as an end-of-stream event. Alternatively, NNG-Stream can be configured to send ZeroMQ push data, and to send empty frames as sentinal values on stream end.
@@ 3.4. Experiment Management with E-Log, ARP and Airflow
The ARP is flexible, and can start different types of workflows tied to run events. The two most common cases are simple shell scripts that submit jobs to a batch queue, or sophisitcated worflows managed by an Apache Airflow instance. Special Airflow operators have been written to interact with the S3DF cluster via the SLURM task management system, and with the Perlmutter supercomputer at NERSC via the Superfacility API.(Enders and others, 2020) An additional operator that will allow communication with Psi-K (see 3.4) is currently under development, with the goal to allow Airflow to schedule tasks on a higher number of HPC facilities using different task scheduling systems.
Since the E-Log/ARP integration is operational, it is simple to enable an Airflow operator to manage an LCLStreamer-API transfer on experiment start/stop. This will start the data streaming pipeline as soon as a data collection run is started. Then compute-intensive analysis jobs will run remotely while the experiment is taking place.
@@ 3.5. HPC Job Management with Psi-K
The Psik and Psik-API projects provide a web-enabled interface for batch queuing systems on local machines and HPC centers. As described above, Psi-k is organized around a file layout where a top-level directory stores a collection of jobs. Documents within a job are rigidly structured. Files contained within each job’s folder (jobs/JobID) include: the JobSpec, a record of job state changes, job logs, and working directory files (as a user-managed collection).
@@
This structure combines ideas from the ExaWorks PSI/J project(Hategan-Marandiuc and others, 2023) with traditional UNIX and cloud-based queuing systems. Note that PSI/J is primarily a script templating and shell-based polling engine, not designed around an asynchronous, web-API model. Hence, it was not suitable for this work.
Jobs are queued by a POST operation, sending a JobSpec json to the jobs path. The server responds with either a validation error or a new JobID. Issuing a GET or a DELETE to jobs/JobID then gets information about the job or cancels a queued or running job. The result is is a single-document method for launching HPC jobs.
@@
Since each JobSpec names its backend, this extra information does not need to be a part of the API path (differing from NERSC’s Superfacility API(Enders and others, 2020)).
@@
Similar to Globus Compute,(Liu and others, 2021; Ananthakrishnan and others, 2024) the API listener can be run by individual users on a compute cluster’s login node. Alternatively, on Oak Ridge’s Frontier system, we run it on a Kubernetes pod with network access to the HPC cluster’s job scheduler and filesystem.(Papadimitriou et al., 2020) Oak Ridge is also implementing S3M API.(Skluzacek et al., 2025) As a client application, Psi-k can send jobs to this API, as well as to the NERSC Superfacility API(Enders and others, 2020) and Globus compute.(Ananthakrishnan and others, 2024) Differing from Globus compute, communications pass directly from the user to the Psik-API, bypasing a cloud services.
@@ 3.6. Mutual authentication with Certified
A companion signer package is designed to issue signed user certificates on a UNIX/Linux system. It can be deployed with essentially zero configuratoin by a typical HPC facility on any login node. Similar to MUNGE (the scheme for issuing user authentication credentials for launching jobs within the SLURM job scheduler) it provides assertions about the user’s login name. In detail, it takes a certificate or a certificate signing request from a user, reads only the user’s public key, and issues the user a certificate linking their public key to their UNIX login name on that system. The user’s login name is determined by asking the kernel for the peer’s SO_PEERCRED information. Differing from MUNGE, it issues signed public keys. There is no danger of exposing these keys, since they can only be used by the original user who posesses the corresponding private key. Certified and signer never send the private key off of the user’s device.
@@
4. Results
This section provides details on the use of LCLStream to address each of the application areas above. For all of the areas, the main result to note was that the framework successfully re-formatted LCLS data for remote consumption and simultaneously transmitted the data over the network to a compute job running on OLCF. The round-trip network latency measured between S3DF’s data transfer nodes and OLCF was consistently around 33-36 milliseconds, only about 50% longer than it takes light to travel the same distance.
All setups made use of parallel producers and parallel consumers. In most cases the bottleneck (determining maximum data throughput) was the data read/formatting speed at S3DF, around 1-3 GB/sec. In other cases, the cache was the rate-limiting step, at around 3-4 GB/sec. Running parallel caches could address the latter, but was not tested in this work.
4.1. Masked X-ray Image Autoencoder (MAXIE) and PeakNet
LCLS operates a variety of different detector types. Thus, the most complicated step in this workflow is curating the input data for training. We produced a series of LCLStreamer input files targeting various experiment/run IDs containing epix10k2M images from the Psana1AssembledAreaDetector source. We also included the photon wavelength read from psana’s SIOC:SYS0:ML00:AO192 as annotation. We opted to implement a special-purpose “PeaknetPreprocessingPipeline” to center and pad images to consistent sizes on the data generation side.
In this application, the LCLStreamer input files replaced a custom data extraction code. The result ended up being more readable and helped with documentation and re-use of the curation steps. Two primary difficulties were noted. First, at this stage of development, it is necessary to manually start the API on S3DF’s data transfer node. Second, we needed to implement our own client-side caching mechanism to prevent re-downloading data. This is significant for this application, since ML training makes many passes over its input. Still, the ability to download needed inputs on-demand decreased the data-gathering lag time between getting an HPC allocation and being able to run scientific work. It also eliminated stale data on the HPC side as a potential source of errors. These results suggest that we need robust, flexible methods like LCLStreamer for pre-processing our data as we work to improve its quality.
4.2. TMO-prefex
The TMO-prefex application was an early version of LCLStreamer. It was used during the commissioning of the TMO end-station just after its upgrade to LCLS-II’s higher data rate. We used Psik-API to run a job computing electron angular and time-of-flight correlation plots on OLCF ACE. Then we manually started MPI-parallel TMO-prefex data producers pushing data through NNG-Stream on S3DF. We synchronized work by interacting through the control room’s live Zoom channel. Due to compression at the source (time-windowed signal thresholding), we did not have a bandwidth limitation. The major source of startup-latency was the program start-up time for parallel data producers on S3DF.
4.3. CrystFEL
CrystFEL has added the capability to work on live data streamed from DECTRIS-type detectors. The data format is specified as part of their Simplon-API, and consists of both control packets and data packets. We implemented this by adding the SimplonBinarySerializer type as an alternative to HDF5 serialization. This serializer inserts the appropriate control messages into the output stream.
During a beamtime at the Macromolecular Femtosecond Crystallography (MFX) endstation at LCLS, the LCLStreamer component was used to send data from one of the experimental runs over to the testbed HPC cluster at the Oak Ridge Leadership Computing Facility, where CrystFEL had been setup to receive and process the transferred data. The latency between data collection and processing with CrystFEL has been shown to be within the range of 15-25 seconds, allowing useful feedback information within a time range sufficient for researchers to manually steer the experiment in the desired direction.
This application especially demonstrated the benefits obtained from standards combined with a modular, interoperable software design. We did not need to create a bespoke data pipeline to generate this message format specifically, nor did CrystFEL have to add a facility-specific data ingestion framework. Instead, we implemented only the specific data format, named it after the standard, and re-used the rest of the facility and user community software pipelines.
@@ 5. Discussion
Data streaming is an emerging technological area within DOE’s IRI (Integrated Research Infrastructure). Our work follows the pattern of the GRETA/Deleria project,(Cromaz et al., 2021) but uses published, open-source, and modular software components targeted to the field of X-ray image analysis. Compared to the edge-to-exascale framework,(Yin et al., 2024) our work focuses on well-defined intermediate data formats and repeatable, modular processing steps. We also target GB/sec data rates arising from high-frequency X-ray sources, as opposed to somewhat lower data rates from synchrotrons or MB/sec data rates from neutron detectors and automated laboratories. The latter types of data flows can still be handled by pub/sub message transport models, as demonstrated by INTERSECT(Thakur et al., 2022) or BlueSky(Wijesinghe et al., 2024) projects. Both the latter frameworks have the somewhat larger goal of controlling online experiments. We do not seek to replace those frameworks, but rather to provide a modular data request and transport mechanism.
Central to this project’s success has been a continued, close collaboration between the DOE-IRI pathfinder projects, LCLS, S3DF, and leadership compute facilities. The design boundaries between the components we eventually arrived at were shaped by sharing ideas and preliminary implementations of workflows among all these groups. The team-developed microservices described here work in concert to accomplish streaming. Continued work on the IRI progam promises even greater integration with facility-deployed infrastructure that will continue to improve their re-usability across sites and reliability for production work.
Metrics for measuring the pace of improvement include: 1) eliminating manual intervention steps during install, startup, and execution (such as starting servers and creating ssh tunnels) 2) reducing site-specific adaptations needed to interoperate across different experiment and compute facility API-s, 3) standardizing security technologies (such as mutual TLS authentication) needed to connect network sockets between data producers and consumers, 4) solving the ”co-scheduling” problem for simultaneous commitment of beam time, network bandwidth, and compute resources, and 5) providing unencumbered access to fast, open-source data movement services (e.g. XRootD or FTS3) for managing site-to-site file copies.
@@ 6. Conclusion
This project has a unique position in relation to other efforts within the DOE Integrated Research Infrastructure (IRI) landscape.(Rogers, 2024) We have chosen to focus on small services that collaborate to provide a high-rate data channel. This avoids dealing with the full complexity of the experimental orchestration layer that Bluesky,(Wijesinghe et al., 2024) INTERSECT,(Brim et al., 2024) and GRETA/DELERIA’s(Cromaz et al., 2021) Janus framework do. At the same time, the certified, psik, and NNG-Stream packages have given general solutions to microservice communication, HPC job submission, and high-rate data buffering. They are targeted to simple, drop-in operations on HPC clusters, and avoid a centralized, cloud controller.
@@
Acknowledgements.
We thank Daniel Pelfrey, Ross Miller, Jordan Webb, Jordan Brown, Carl Bai, and Matthew Ezell for configuring high-speed network access from the Defiant system on OLCF’s ACE testbed and from compute nodes on OLCF Frontier. This research used resources of the SLAC National Accelerator Laboratory and the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. Use of the Linac Coherent Light Source (LCLS), SLAC National Accelerator Laboratory, is supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under Contract No. DE-AC02-76SF00515.